OpenStudy (anonymous):

Quick derivative question? I dont understand what d^2y/dx^2 means. I thought it was just the derivitive of the derivitive but this is wrong. x = 6t + ln(t) y= 3t-ln(t)

OpenStudy (anonymous):

the x and y is what i need to get help with.

OpenStudy (e.mccormick):

Typically, \[f''(x)=(f'(x))'=y''=\frac{d^2y}{dx^2}=\frac{d}{dx} ( \frac{dy}{dx} ) \]

OpenStudy (anonymous):

thanks! i cant believe i didnt understand that..

OpenStudy (e.mccormick):

Now, what you put in below the derivative part was something with a t, so this may be something completely different.

OpenStudy (e.mccormick):

And yah, the derivative of the derivative is one way to say second derivative.